There are a total of 10 test stations. When a circuit board fails, the test station will automatically record the lifetime for the failed board and then immediately switch to a replacement board. The test was stopped at 1200 hours. At that time, 19 circuit boards had failed and there was an operating circuit board in every station. Assuming the boards have exponentially-distributed lifetimes, estimate the mean lifetime for this type of board. State your answer rounded to the nearest hour.
Answer #1

Solutions Here we estimate the mean lifetime for this type of board, from the given data Total test stations = 10 test was stworking from x - log (12) - 0.6931 mean working time Cut) X- T; 7=1200 Ha LX1200 $ 437 1200 437 - 2.746 F(x) = 1/2 x= - 437 l